# Semi-parametric bulk and tail regression using spline-based neural networks

### v1

@@ 3.4 Inference
Inference for xSPQR follows similarly to SPQR, as described in Section 2.2. We use Adam, with default hyper-parameters, to train the MLP that defines the parameter vector $(\xi(\mathbf{x}),w_{1}(\mathbf{x}),\dots,w_{K}(\mathbf{x}))$, where the loss function is the negative log-likelihood associated with the conditional bGP density in (7). We note that the hyper-parameters of the blending function $p(y)$ in (4) are fixed during training of the model. Prior to model fitting, observations $\{(y_{i},\mathbf{x}_{i})\}_{i=1}^{n}$ are split into training and test sets, $\{(y_{i},\mathbf{x}_{i})\}_{\rm train}$ and $\{(y_{i},\mathbf{x}_{i})\}_{\rm test}$, respectively. We assume our observations $y_{i}$ satisfy $y_{i}>0,i=1,\dots,n$; however, as the SPQR density is supported on the unit interval, and training response observations $\{y_{i}\}_{\rm train}$ may not fall in $[0,1]$, we must first standardise $\{y_{i}\}_{\rm train}$ to the unit interval via a min-max transformation; this is based on the training data alone. We use the same values to standardise $\{y_{i}\}_{\rm test}$, however, there are no guarantees that all observations of the resulting standardised test response data will satisfy $y_{i}<1$. Although this is not an issue for xSPQR, any such test observation cannot be handled by SPQR without re-training of the model, and all associated probability and density estimates will degenerate to one and zero, respectively; examples of this phenomena are illustrated in the application, Section 5.
@@
The xSPQR model is trained using the R interface to keras (Allaire and Chollet, , 2021). Before training, we further hold out $20\%$ of the training data for validation during optimisation of the MLP. To mitigate overfitting, our training scheme uses checkpoints and early-stopping (Prechelt, , 2002). At each iteration of the optimisation algorithm, we evaluate the negative log-likelihood on the validation data, and then determine the “best fitting” model (over all iterations) as that which minimises the validation loss function. Early-stopping halts the optimisation scheme early if the validation loss has not decreased in the last $\Delta\in\mathbb{N}$ iterations (we set $\Delta=25$ throughout). We also add $L_{1}$ penalties to the final activation layer for the shape parameter, $\xi(\mathbf{x})$, which provides further model regularisation; see, e.g., Goodfellow et al., (2016).
@@
Courtesy of the modular nature of the xSPQR density, we can use pre-training (see Goodfellow et al., (2016), Ch.8) to increase the computational efficiency of the inference procedure and mitigate numerical instability during training. When fitting the xSPQR model, we first fit the corresponding SPQR model, i.e., with $p_{a}=p_{b}\rightarrow 1$. We then use the parameters of the estimated neural network as initial estimates when training xSPQR. We find that this procedure works well in practice.
@@
While we presented the xSPQR method in Section 3 for a real-valued shape parameter, $\xi(\mathbf{x})\in\mathbb{R}$, there may be computational and practical benefits to constraining $\xi(\mathbf{x})$ to be non-negative. When the shape parameter $\xi(\mathbf{x})$ is negative, the resulting GP distribution has a finite upper endpoint, which can make validation and testing of GP regression models computationally troublesome. If test or validation response data are observed to exceed their predicted upper-endpoint, then the negative log-likelihood function will evaluate to a non-finite value and, as a consequence, the loss surface is highly non-regular; iterative stochastic gradient descent methods (e.g., Adam) may then struggle to find global maxima. For a discussion of this issue, particularly with respect to deep GP regression models, see Richards and Huser, (2022, 2024). For simplicity, we hereafter constrain $\xi(\mathbf{x})\in(0,1)$ using the sigmoid activation function in the final layer of the MLP; the choice to constrain $\xi(\mathbf{x})<1$ is judicious, and is motivated primarily by the typical values of $\xi(\mathbf{x})$ found in applications of GP regression models.
3.5 Relative variable importance for xSPQR
@@
As with many applications of density regression models, we seek to assess the relative importance of covariates on determining the conditional distribution. In the xSPQR framework, the parameters of the conditional density comprise those that influence the bulk (i.e., $\mathcal{W}(\mathbf{x})$) and those that are specific to the tails (i.e., $\xi(\mathbf{x})$). Thus, xSPQR permits the covariates to separately affect the body and the tail of the conditional distribution; as a result of this, we may expect the relative importance of covariates to change between the bulk and tails.
@@
Following the discussion in Section 2.3, we assess relative importance in the bulk by estimating the VI scores for a sequence of $\tau$ values; recall that, for $\tau<p_{a}$, the conditional distribution function for the xSPQR model is exactly that of the usual SPQR framework. When considering relative importance in the tails, we have two options: i) considering VI scores for $\tau$ close to one or ii) directly interpreting relative importance for the shape parameter, $\xi(\mathbf{x})$; the latter can be achieved similarly to the conditional quantile function (for fixed $\tau$), by setting $g(\mathbf{x})$ in (2.3) to be $\xi(\mathbf{x})$.
@@ 4.1 Overview
Here we provide a simulation study to showcase the benefits of using xSPQR over SPQR when extrapolating into the tails. Covariates $\mathbf{X}_{i},i=1,\dots,3$, are drawn independently from a Unif(0,1) distribution. We then draw our response $Y\mid(\mathbf{X}=\mathbf{x})$ from a log-normal distribution with density function $f(y|\mathbf{x})=(\sqrt{2\pi}\sigma(\mathbf{x})y)^{-1}\exp\{-(\log y-\mu(\mathbf{x}))^{2}/(2\sigma^{2}(\mathbf{x}))\}$ for $y>0,\sigma(\mathbf{x})>0,$ and $\mu(\mathbf{x})\in\mathbb{R}$. The parameters $\mu(\mathbf{x})$ and $\sigma(\mathbf{x})$ are dependent on $\mathbf{x},$ and we set $\mu(\mathbf{x})=5(1-1/(1+\exp(-(1-5x_{1}x_{2}))))$ and $\sigma(\mathbf{x})=1/(1+\exp(-(1-5x_{1}x_{2})))$; hence only $X_{1}$ and $X_{2}$ act on $Y$.
@@
$\mbox{IWD}=\int_{\mathcal{X}}\int^{1}_{0}|Q(y|\mathbf{x})-\hat{Q}(y|\mathbf{x})|\mathrm{d}\mathbf{x},$
where $Q(y|\mathbf{x})$ denotes the true conditional quantile function (and $\hat{Q}(y|\mathbf{x})$ denotes its estimate), and where $\mathcal{X}$ denotes the sample space for the covariate vector $\mathbf{X}$ (in the above case, $\mathcal{X}=[0,1]^{3}$). We estimate the IWD out-of-sample by generating a test set of 5000 covariate vectors and performing Monte-Carlo integration. We also consider a tail-calibrated version of the IWD, denoted by tIWD, which is constructed by replacing the limits of the inner integral of (4.1) with $[0.999,1]$.
We perform 100 experiments; for each, we generate $n\in\{1000,10000\}$ samples from the aforementioned log-normal regression model, and fit the original SPQR and our xSPQR models. We consider a coarse grid of hyper-parameters for the two models, which was motivated by unreported initial experiments. For the $M$-spline basis and underlying MLP (of both models), we consider $K\in\{15,25\}$ basis functions, a $2$-layered neural network with $n_{h}\in\{16,32\}$ nodes per layer, and two activation functions $\bm{\phi}(\cdot)$: sigmoid and ReLU. For the hyper-parameters specific to xSPQR, we consider $p_{a}\in\{0.75,0.9,0.925\}$, $p_{b}\in\{0.95,0.99,0.999\},$ and $c_{1}\in\{5,25\}$. The $\xi$-specific activation function is taken to be a sigmoid activation, scaled such that $\xi(\mathbf{x})\in(0,0.5)$.
@@ 4.2 Results
Table 1: Results of the simulation study. The median of the IWD and tIWD estimates (alongside 50% confidence intervals; in brackets) are reported for the original SPQR and new xSPQR models, with values for the latter in bold. The hyper-parameters of the xSPQR, $(p_{a},p_{b},c_{1})$, are optimised for each row by minimising the tIWD, and are reported in the final column.
@@
Table 1 provides the results of the simulation study. For brevity, we present here only the results when $\bm{\phi}(\cdot)$ in (2.2) is the sigmoid activation; this provided consistently better results compared to the ReLU activation function. We observe that xSPQR consistently provides lower IWD and tIWD estimate than the competing SPQR method, which illustrates the efficacy of xSPQR for modelling heavy-tailed conditional densities. As one might expect, the estimation accuracy tends to increase with the number of knots and the number of nodes in the underlying MLP. For the optimal hyper-parameters, we find a lower $c_{1}$ is preferred in a low sample setting, whilst a higher $c_{1}$ is preferred when more data are available for inference. As discussed alongside Figure 1, a higher value of $c_{1}$ puts more weight on the more flexible SPQR distribution during blending; this may be preferable when you have more data available to infer this component of the model. For the blending interval, we find consistent agreement, when $n=10000$, that the preferred blending quantile levels are $p_{a}=0.75$ and $p_{b}=0.99$; this provides a relatively wide interval for blending of the two constituent distribution functions, with the exact GP upper-tails beginning at $b$. Although Table 1 provides some intuition as to appropriate choices of hyper-parameters for xSPQR, it is by no means conclusive. We advocate that, in practice, a grid-search over hyper-parameters is performed and the quality of xSPQR fits are compared on hold-out data.
@@
Figure 3: Simulation study: estimates of density (top) and log-survival (bottom) functions for three test covariate vectors. The red dashed lines give the true functions; the black and green curves are the corresponding estimates from SPQR and xSPQR, respectively. The values of $a$ and $b$ (from the xSPQR fit) are denoted by the orange horizontal lines.
@@
Figure 4: Simulation study: estimates of the variable importance scores $\mbox{VI}_{j}(\hat{Q}(\tau|\mathbf{x}))$ for the covariates $X_{j},j=1,2,3,$ as a function of the quantile level $\tau$.
@@
Figure 3 compares out-of-sample density and distribution estimation using SPQR and xSPQR. In this case, we take model fits to $n=10000$ samples, using the optimal hyper-parameters from the bottom row of Table 1. As Figure 3 illustrates, xSPQR is much better able to capture the upper-tails of the true log-normal distribution function. Moreover, SPQR is unable to provide any useful estimates of the conditional distribution function for $y>1$. Figure 4 also provides estimates of the variable importance scores from the aforementioned fitted xSPQR model. Specifcally, we estimate $\mbox{VI}_{j}(\hat{Q}(\tau|\mathbf{x}))$ where the function of interest $g(\cdot)$, defined in 5, is taken to be the estimated conditional $\tau$-quantile function. The estimates are able to capture the true underlying covariate importance, specifically that $X_{3}$ does not have any effect on the conditional distribution at any quantile level $\tau$, whilst $X_{1}$ and $X_{2}$ have an equal effect for all values of $\tau$. The $\xi$-specific variable importance scores (see Section 3.5) are in agreement; their estimates are $0.0156,0.0234,0.0009$ for $X_{1}$, $X_{2}$, and $X_{3}$, respectively.
@@
[…]
@@
Following Lawler and Shaby, (2024), we split the data into training and testing sets of 6416 and 3344 fires, respectively. The split is not random, as the test set consists of the first and last five years of the data: 1184 fires from 1990 to 1994, inclusive, and 2160 fires from 2016 to 2020, inclusive. Additionally, we uniformly-at-random select $20\%$ of the training data for validation during training of the xSPQR model; see Section 3.4. As illustrated by Richards and Huser, (2022) and Lawler and Shaby, (2024), U.S. wildfire burnt areas are extremely heavy-tailed ($\xi>0.5$), which can cause numerical difficulties during training of deep extreme quantile regression models (Richards and Huser, , 2022). While Lawler and Shaby, (2024) employ a log-transform of the data to mitigate this issue, we instead follow Richards and Huser, (2022) and model the square-root of the burnt areas as our response variable, $Y$; the square-root transformation better preserves the tail behavior of the response compared to the log-transform, and can be interpreted as the diameter of an affected region. The wildfires are assumed to be independent of each other; while a naive assumption, it is nevertheless aligned with our specific goal of testing the appropriateness of the xSPQR framework for density regression.
@@
We fit and compare both the original SPQR and the EVT-compliant xSPQR models to the data. When deciding on the optimal hyper-parameters for the SPQR/xSPQR model, we perform a coarse grid search over hyper-parameters to choose their optimal values; we use the same MLP architecture for both frameworks. The candidate models are compared using the negative log-likelihood (evaluated on the test data) and by visual examination of a model diagnostic, the pooled QQ plot which is to be discussed. The optimal neural network comprising our SPQR/xSPQR model has $H=2$ hidden layers, each with $n_{h}=12$ nodes, and with $K=25$ basis functions describing the body of the distribution. An exploratory data analysis, whereby we fit stationary GP models to the response data, reveals that the response burnt area data are in fact heavy-tailed, with the vast majority of fits returning estimates of $\xi$ exceeding zero. As such, we constrain $\xi(\mathbf{x})>0$ for xSPQR by using the logistic activation function. For the blending interval, we set $p_{a}$ and $p_{b}$ to be $0.9$ and $0.999$, respectively. Finally, $c_{1}$ and $c_{2}$ are set to be $25$ and $5$, respectively, with these two values being carried over from our numerical studies.
@@
The covariates for both the training and test sets are normalised by subtracting and dividing by the marginal mean and standard deviation vectors, respectively, of the training covariate data. Similarly, the response data are scaled to $[0,1]$ using the min-max transformation, as described in Section 3.4. As a consequence, there are two response values in the scaled test data which exceed one, corresponding to the two largest wildfires across the entire dataset, which occurred in 2017 and 2020. As explained in Section 3.4, regular SPQR can not provide any usable estimates for these particularly extreme test values, but this is not an issue for xSPQR.
@@
Figure 5: Goodness-of-fit diagnostics for the SPQR and xSPQR models. Pooled PP (top) and QQ (bottom) plots for the fitted SPQR (left) and xSPQR (right) models, evaluated on the test data. The red points in the right column denote the two most extreme test values, which cannot be modelled using SPQR and thus do not appear on the left column.
@@
Figure 5 provides visual goodness-of-fit diagnostics for the fitted SPQR and xSPQR models. We create out-of-sample pooled PP and QQ plots, whereby the fitted conditional models are used to transform the test data onto standard (unconditional) margins. The top row of PP plots are on uniform margins which showcase goodness-of-fit for the entire distribution, while the bottom row QQ plots are on exponential margins, to highlight the fits in the upper-tails (see, e.g., Heffernan and Tawn, , 2001); deviation of points from the diagonal suggest poor fits. Both SPQR and xSPQR provide fantastic fits for the entire range of the data, however, xSPQR gives markedly better fits in the upper-tails, as the model is better designed to handle extreme events. Moreover, xSPQR is capable of predicting well the probability of the two most extreme burnt areas in the test data; SPQR is unable to provide any usable estimates for these events.
@@
Figure 6: Estimates of the variable importance scores $\mbox{VI}_{j}(\hat{Q}(\tau|\mathbf{x}))$ as a function of the quantile level $\tau$. Different colours correspond to different covariates, while the dashed and solid lines correspond to estimates from SPQR and xSPQR, respectively. The left and right panels correspond to the two test periods: 1990–1994 and 2016–2020, respectively; the central panel corresponds to the training period, 1995–2015.

### v2

@@ 3.4 Inference
Inference for SPQRx follows similarly to SPQR, as described in Section 2.2. We use Adam, with default hyper-parameters, to train the MLP that defines the parameter vector $(\xi(\mathbf{x}),w_{1}(\mathbf{x}),\dots,w_{K}(\mathbf{x}))$, where the loss function is the negative log-likelihood associated with the conditional bGP density in (7) and a penalty term to encourage positivity of the conditional density function in the blending intervals (see discussion in Section 3.2 and Appendix B). We note that the hyper-parameters of the blending function $p(y)$ in (4) are fixed during training of the model. Prior to model fitting, observations $\{(y_{i},\mathbf{x}_{i})\}_{i=1}^{n}$ are split into training and test sets, $\{(y_{i},\mathbf{x}_{i})\}_{\rm train}$ and $\{(y_{i},\mathbf{x}_{i})\}_{\rm test}$, respectively. We assume our observations $y_{i}$ satisfy $y_{i}>0,i=1,\dots,n$. However, as the SPQR density is supported on the unit interval, and training response observations $\{y_{i}\}_{\rm train}$ may not fall in $[0,1]$, we must first standardise $\{y_{i}\}_{\rm train}$ to the unit interval via a min-max transformation; this is based on the training data alone. We use the same values to standardise $\{y_{i}\}_{\rm test}$, however, there are no guarantees that all observations of the resulting standardised test response data will satisfy $y_{i}<1$. Although this is not an issue for SPQRx, any such test observation cannot be handled by SPQR without re-training of the model, and all associated probability and density estimates will degenerate to one and zero, respectively. Examples of this phenomena are illustrated in the application, Section 5.
@@
The SPQRx model is trained using the R interface to keras (Allaire and Chollet, 2021). Before training, we further hold out $20\%$ of the training data for validation during optimisation of the MLP. To mitigate overfitting, our training scheme uses checkpoints and early-stopping (Prechelt, 2002). At each iteration of the optimisation algorithm, we evaluate the negative log-likelihood on the validation data, and then determine the “best fitting” model (over all iterations) as that which minimises the validation loss function. Early-stopping halts the optimisation scheme early if the validation loss has not decreased in the last $\Delta\in\mathbb{N}$ iterations (we set $\Delta=25$ throughout). We also add $L_{1}$ penalties to the final activation layer for the shape parameter, $\xi(\mathbf{x})$, which provides further model regularisation; see, e.g., Goodfellow et al. (2016).
@@
Courtesy of the modular nature of the SPQRx density, we can use pre-training (see Goodfellow et al. (2016), Ch.8) to increase the computational efficiency of the inference procedure and mitigate numerical instability during training. When fitting the SPQRx model, we first fit the corresponding SPQR model, i.e., with $p_{a}=p_{b}\rightarrow 1$. We then use the parameters of the estimated neural network as initial estimates when training SPQRx. We find that this procedure works well in practice.
@@
While we presented the SPQRx method in Section 3 for a real-valued shape parameter, $\xi(\mathbf{x})\in\mathbb{R}$, there may be computational and practical benefits to constraining $\xi(\mathbf{x})$ to be non-negative. When the shape parameter $\xi(\mathbf{x})$ is negative, the resulting GP distribution has a finite upper endpoint, which can make validation and testing of GP regression models computationally troublesome. If test or validation response data are observed to exceed their predicted upper-endpoint, then the negative log-likelihood function will evaluate to a non-finite value. As a consequence, the loss surface is highly non-regular, and iterative stochastic gradient descent methods (e.g., Adam) may then struggle to find global maxima. For a discussion of this issue, particularly with respect to deep GP regression models, see Richards and Huser (2026b, a). If negative values of $\xi(\mathbf{x})$ are desirable, the training scheme of Mackay et al. (2025) can be employed. They advocated initialising the model with $\xi(\mathbf{x})$ strictly positive for all $\mathbf{x}$. The loss function is thus guaranteed to be finite at the onset of training. Then, if the optimisation algorithm produces non-finite loss values during training, training restarts from the last iteration (with a finite loss) and the learning rate is reduced. This procedure is repeated for a sequence of decreasing learning rate values.
3.5 Relative variable importance for SPQRx
@@
As with many applications of density regression models, we seek to assess the relative importance of covariates on determining the conditional distribution. In the SPQRx framework, the parameters of the conditional density comprise those that influence the bulk (i.e., $\mathcal{W}(\mathbf{x})$) and those that are specific to the tails (i.e., $\xi(\mathbf{x})$). Thus, SPQRx permits the covariates to separately affect the body and the tail of the conditional distribution. As a result of this, we may expect the relative importance of covariates to change between the bulk and tails.
@@
Following the discussion in Section 2.3, we assess relative importance in the bulk by estimating the VI scores for a sequence of $\tau$ values; recall that, for $\tau<p_{a}$, the conditional distribution function for the SPQRx model is exactly that of the usual SPQR framework. When considering relative importance in the tails, we have two options: i) considering VI scores for $\tau$ close to one or ii) directly interpreting relative importance for the shape parameter, $\xi(\mathbf{x})$. The latter can be achieved similarly to the conditional quantile function (for fixed $\tau$), by setting $g(\mathbf{x})$ in (2.3) to be $\xi(\mathbf{x})$.
@@ 4.1 Overview
Here we provide a simulation study to showcase the benefits of using SPQRx over SPQR and competing deep quantile regression methods, when extrapolating into the tails. Covariates $\mathbf{X}_{i},i=1,\dots,3$, are drawn independently from a Unif(0,1) distribution. We then draw our response $Y\mid(\mathbf{X}=\mathbf{x})$ from a log-normal distribution with density function $f(y|\mathbf{x})=(\sqrt{2\pi}\sigma(\mathbf{x})y)^{-1}\exp\{-(\log y-\mu(\mathbf{x}))^{2}/(2\sigma^{2}(\mathbf{x}))\}$ for $y>0,\sigma(\mathbf{x})>0,$ and $\mu(\mathbf{x})\in\mathbb{R}$. The parameters $\mu(\mathbf{x})$ and $\sigma(\mathbf{x})$ are dependent on $\mathbf{x},$ and we set $\mu(\mathbf{x})=5(1-1/(1+\exp(-(1-5x_{1}x_{2}))))$ and $\sigma(\mathbf{x})=1/(1+\exp(-(1-5x_{1}x_{2})))$; hence only $X_{1}$ and $X_{2}$ act on $Y$.
@@
$\mbox{IWD}=\int_{\mathcal{X}}\int^{1}_{0}\left|Q(y|\mathbf{x})-\hat{Q}(y|\mathbf{x})\right|\mathrm{d}\mathbf{x},$
where $Q(y|\mathbf{x})$ denotes the true conditional quantile function (and $\hat{Q}(y|\mathbf{x})$ denotes its estimate), and where $\mathcal{X}$ denotes the sample space for the covariate vector $\mathbf{X}$ (in the above case, $\mathcal{X}=[0,1]^{3}$). We estimate the out-of-sample IWD by generating a test set of 5000 covariate vectors and performing Monte-Carlo integration. We also consider a tail-calibrated version of the IWD, denoted by tIWD, which is constructed by replacing the limits of the inner integral of (4.1) with $[0.999,1]$.
We perform 100 experiments. For each, we generate $n\in\{1000,10000\}$ samples from the aforementioned log-normal regression model, and fit the original SPQR model, the SPQRx model, and a deep GP regression model (Wilson et al., 2022; Richards et al., 2023), using the parameterisation proposed by Pasche and Engelke (2024). The deep GP model involves a two-step estimation procedure: first, a deep quantile regression model is used to estimate a non-stationary exceedance threshold, corresponding to the conditional $p_{a}$-quantile (for $p_{a}$ close to one). We then fit a non-stationary deep GP model to excesses above this threshold. The intermediate exceedance quantile and the deep GP parameters are represented as MLPs with the same architecture (i.e., number of layers and width per layer). Note that the deep GP regression model can only describe the conditional quantile function, $Q(y|\mathbf{x}),$ for $y\geq Q(p_{a}|\mathbf{x})$. Thus, we cannot evaluate the IWD for this model.
For all three models, we consider a coarse grid of hyper-parameters which is motivated by unreported initial experiments. For the $M$-spline basis of the SPQR density, we consider $K\in\{15,25\}$ basis functions. For fairness of comparison, we consider the same underlying architectures for all aforementioned MLPs (SPQR, SPQRx, and the two that comprise the deep GP model): a $2$-layered neural network with $n_{h}\in\{16,32\}$ nodes per layer, and two activation functions $\bm{\phi}(\cdot)$: sigmoid and ReLU. For the hyper-parameters specific to SPQRx, we consider $p_{a}\in\{0.75,0.9\}$, $p_{b}\in\{0.99,0.999\},$ and $c_{1}\in\{5,25\}$. For the deep GP regression model, we consider $p_{a}\in\{0.75,0.9,0.925\}$. For both SPQRx and the deep GP model, the $\xi$-specific activation function is taken to be the hyperbolic tangent function, scaled and shifted to ensure that $\xi(\mathbf{x})\in(-0.5,0.7)$ (see Pasche and Engelke, 2024). In both cases, we use the iterative training scheme described in Section 3.4 to accommodate the negative shape values.
@@ 4.2 Results
Table 1: Results of the simulation study. The median of the IWD and tIWD estimates (alongside 50% confidence intervals; in brackets) are reported for the original SPQR, new SPQRx model (in bold), and the deep GP regression model (underlined). Note that the IWD is undefined for the deep GP model and the hyperparameter $K$ is not used in its construction. The hyper-parameters of the SPQRx, $(p_{a},p_{b},c_{1})$, are optimised for each row by minimising the tIWD, and are reported in the final column.
@@
Table 1 provides the results of the simulation study. For brevity, we present here only the results when $\bm{\phi}(\cdot)$ in (2.2) is the sigmoid activation, which provides consistently better results compared to the ReLU activation function. We also present only the optimal results for the deep GP regression model. We observe that SPQRx generally provides lower IWD and tIWD estimate than the competing SPQR method, which illustrates the efficacy of SPQRx for modelling heavy-tailed conditional densities. The SPQRx model consistently outperforms the deep GP regression model, and further benefits from a full description of the conditional distribution not offered by peaks-over-threshold models. As one might expect, the estimation accuracy tends to increase with the number of knots and the number of nodes in the underlying MLP. For the optimal hyper-parameters of SPQRx, we find a lower $c_{1}$ is preferred in a low sample setting, whilst a higher $c_{1}$ is preferred when more data are available for inference. As discussed alongside Figure 1, a higher value of $c_{1}$ puts more weight on the more flexible SPQR distribution during blending. This may be preferable when you have more data available to infer this component of the model. For the blending interval, we find consistent agreement, when $n=1000$, that the preferred blending quantile levels are $p_{a}=0.9$ and $p_{b}=0.99$. This provides a relatively wide interval for blending of the two constituent distribution functions, with the exact GP upper-tails beginning at $b$. Although Table 1 provides some intuition as to appropriate choices of hyper-parameters for SPQRx, it is by no means conclusive. We advocate that, in practice, a grid-search over hyper-parameters is performed and the quality of SPQRx fits are compared on hold-out data.
@@
Figure 3: Simulation study: estimates of density (top) and log-survival (bottom) functions for three test covariate vectors. The red dashed lines give the true functions. The black and green curves are the corresponding estimates from SPQR and SPQRx, respectively. The values of $a$ and $b$ (from the SPQRx fit) are denoted by the orange horizontal lines.
@@
Figure 4: Simulation study: estimates of the variable importance scores $\mbox{VI}_{j}\left(\hat{Q}(\tau|\mathbf{x})\right)$ for the covariates $X_{j},j=1,2,3,$ as a function of the quantile level $\tau$.
@@
Figure 3 compares out-of-sample density and distribution estimation using SPQR and SPQRx. In this case, we take model fits to $n=10000$ samples, using the optimal hyper-parameters from the bottom row of Table 1. As Figure 3 illustrates, SPQRx is much better able to capture the upper-tails of the true log-normal distribution function. Moreover, SPQR is unable to provide any useful estimates of the conditional distribution function for $y>1$. Figure 4 also provides estimates of the variable importance scores from the aforementioned fitted SPQRx model. Specifically, we estimate $\mbox{VI}_{j}\left(\hat{Q}(\tau|\mathbf{x})\right)$ where the function of interest $g(\cdot)$, defined in 5, is taken to be the estimated conditional $\tau$-quantile function. The estimates are able to capture the true underlying covariate importance, specifically that $X_{3}$ does not have any effect on the conditional distribution at any quantile level $\tau$, whilst $X_{1}$ and $X_{2}$ have an equal effect for all values of $\tau$. The $\xi$-specific variable importance scores (see Section 3.5) are in agreement, with estimates $0.0156,0.0234,0.0009$ for $X_{1}$, $X_{2}$, and $X_{3}$, respectively.
Two additional simulation studies are described in Appendix C. Here we consider the higher-dimensional setting, where $p=20$, and where the heaviness of the conditional tails is allowed to vary with the covariates (i.e., $\xi(\mathbf{x})\neq 0$ for all $\mathbf{x}$). We find that SPQRx outperforms SPQR in both studies. When data are heavy-tailed, SPQRx also outperforms the deep GP model, but performs slightly worse when the data have bounded upper-tails, i.e., $\xi(\mathbf{x})<0$ for all $\mathbf{x}$. In this case, estimation of the SPQRx model can be computationally troublesome (see Section 3.4). In our application to U.S. wildfire burnt areas, the data are demonstrably heavy-tailed.
@@
We demonstrate the flexibility of SPQRx by modelling U.S. wildfire burnt areas, that were originally compiled and analysed by Lawler and Shaby (2024). The data consist of burnt areas (in 1000s of acres) and counts for over 10,000 wildfires across the contiguous U.S., from 1990 to 2020, with only fires exceeding 1000 acres of burnt area included in the dataset. The dataset includes several covariates, including spatio-temporal coordinates, meteorological variables, fire weather indices, and housing density. Lawler and Shaby (2024) studied these data using a joint model for counts and area with spatial random effects, and employed the extended GP distribution for modelling burnt areas. In our work, we focus on modelling the distribution of the continuous burnt areas, conditional on the following 6 covariates: year (fire_yr), total precipitation during the month of the fire (pr_curr; mm) and during the entire previous year (pr_prev; mm), minimum relative humidity (rmin; $\%$-age), maximum temperature (tmax; K), and windspeed (wspd; m/s). The final three variables (rmin, tmax, wspd) are aggregated into their monthly means.
@@
Following Lawler and Shaby (2024), we split the data into training and testing sets of 6416 and 3344 fires, respectively. The split is not random, as the test set consists of the first and last five years of the data: 1184 fires from 1990 to 1994, inclusive, and 2160 fires from 2016 to 2020, inclusive. Additionally, we uniformly-at-random select $20\%$ of the training data for validation during training of the SPQRx model; see Section 3.4. As illustrated by Richards and Huser (2026b) and Lawler and Shaby (2024), U.S. wildfire burnt areas are extremely heavy-tailed ($\xi>0.5$), which can cause numerical difficulties during training of deep extreme quantile regression models (Richards and Huser, 2026b). While Lawler and Shaby (2024) employ a log-transform of the data to mitigate this issue, we instead follow Richards and Huser (2026b) and model the square-root of the burnt areas as our response variable, $Y$. The square-root transformation better preserves the tail behavior of the response compared to the log-transform, and can be interpreted as the diameter of an affected region. The wildfires are assumed to be independent of each other; while a naive assumption, it is nevertheless aligned with our specific goal of testing the appropriateness of the SPQRx framework for density regression.
@@
We fit and compare both the original SPQR and the EVT-compliant SPQRx models to the data. When deciding on the optimal hyper-parameters for the SPQR/SPQRx model, we perform a coarse grid search over hyper-parameters to choose their optimal values. We use the same MLP architecture for both frameworks. The candidate models are compared using the negative log-likelihood (evaluated on the test data) and by visual examination of a model diagnostic, the pooled QQ plot, which is to be discussed. The optimal neural network comprising our SPQR/SPQRx model has $H=2$ hidden layers, each with $n_{h}=12$ nodes, and with $K=25$ basis functions describing the body of the distribution. An exploratory data analysis, whereby we fit stationary GP models to the response data, reveals that the response burnt area data are in fact heavy-tailed, with the vast majority of fits returning estimates of $\xi$ exceeding zero. For simplicity, we thus constrain $\xi(\mathbf{x})>0$ for SPQRx by using the logistic activation function. This avoids the need for the burdensome training scheme described in Section 3.4. For the blending interval, we set $p_{a}$ and $p_{b}$ to be $0.9$ and $0.999$, respectively. Finally, $c_{1}$ and $c_{2}$ are set to be $25$ and $5$, respectively, with these two values being carried over from our numerical studies. Most of these settings are chosen based on an initial unreported coarse grid-search over hyper-parameters. We anticipate that relaxing the assumption of $\xi(\mbox{\bf x})>0$, or choosing $(c_{1},c_{2})$ based on the grid search, could change the ‘optimal’ architecture for our analysis.
@@
The covariates for both the training and test sets are normalised by subtracting and dividing by the marginal mean and standard deviation vectors, respectively, of the training covariate data. Similarly, the response data are scaled to $[0,1]$ using the min-max transformation, as described in Section 3.4. As a consequence, there are two response values in the scaled test data that exceed one, corresponding to the two largest wildfires across the entire dataset, which occurred in 2017 and 2020. As explained in Section 3.4, regular SPQR can not provide any usable estimates for these particularly extreme test values, but this is not an issue for SPQRx. Estimation uncertainty is quantified via 200 bootstrap samples of the training period data. Note that the validation data change with every bootstrap sample.
@@
Figure 5: Goodness-of-fit diagnostics for the SPQR and SPQRx models. Pooled $P$ (top) and QQ (bottom) plots for the fitted SPQR (left) and SPQRx (right) models, evaluated on the test data. The red points in the right column denote the two most extreme test values, which cannot be modelled using SPQR and thus do not appear on the left column. The grey area bounded by the dashed lines represent the $95\%$ bootstrap confidence intervals.
@@
Figure 5 provides visual goodness-of-fit diagnostics for the fitted SPQR and SPQRx models. We create out-of-sample pooled PP and QQ plots, whereby the fitted conditional models are used to transform the test data onto standard (unconditional) margins. The top row of PP plots are on uniform margins which showcase goodness-of-fit for the entire distribution, while the bottom row QQ plots are on exponential margins, to highlight the fits in the upper-tails (see, e.g., Heffernan and Tawn, 2001); deviation of points from the diagonal suggest poor fits. Both SPQR and SPQRx provide fantastic fits for the entire range of the data. However, SPQRx benefits from an asymptotically-justified mechanism for extrapolating into the upper-tails. This is illustrated by the red points in Figure 5, which correspond with the two most extreme burnt areas in the test data. While SPQRx is capable of predicting well these events, SPQR is unable to provide any usable probability statements, as they are outside of the range of the training data (see Section 3.4).
@@
Figure 6: Estimates of the variable importance scores $\mbox{VI}_{j}\left(\hat{Q}(\tau|\mathbf{x})\right)$ as a function of the quantile level $\tau$. Different colours correspond to different covariates, while the dashed and solid lines correspond to estimates from SPQR and SPQRx, respectively. The left and right panels correspond to the two test periods: 1990–1994 and 2016–2020, respectively. The central panel corresponds to the training period, 1995–2015.